OpenStudy (anonymous):

If a moving van and a skate board are both going down hill at 60 mph, which one has the most momentum? (20 points) A. the van B. the skateboard C. they are the same D. not enough information

OpenStudy (anonymous):

@alibaby

OpenStudy (anonymous):

okay please give me a second

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

i though b

OpenStudy (anonymous):

im sorry i dont know this but may be @dan815 could help

OpenStudy (anonymous):

can i give u another question

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

In 1987, Marisa Canofoglia, of Italy, roller-skated at a record-setting speed of 11 m/s. If the magnitude of Canofoglia’s momentum was 6.60 × 10^2 kg•m/s, what was her mass? (20 points) Extra Content p=mv or m=p/v or v=p/m A. 7260 kg B. 660 kg C. 60 kg D. 66 kg

OpenStudy (anonymous):

i was thinking b

OpenStudy (anonymous):

i was thinking the same thing

OpenStudy (anonymous):

i change my mind i think its a

OpenStudy (anonymous):

y do you think it's A

OpenStudy (anonymous):

bc i worked out the equation it gave me and then i multiplyed it with 11

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

um im not sure please wait i still think it's b but i dont want u too get the answer wrong @nincompoop @pooja195

OpenStudy (anonymous):

alright

OpenStudy (danjs):

read up on momentum = mass x velocity

OpenStudy (carolyn.x):

A, because momentum refers to the product of mass and velocity, since van has a higher mass, and both travel at 60mph so can has higher momentum

OpenStudy (carolyn.x):

Van*